# Life on a closed timelike curve

### v1

@@ Abstract
We study the internal dynamics of a hypothetical spaceship traveling on a close timelike curve in a Gödel-type universe. We choose the curve so that the generator of evolution in proper time is the angular momentum. Using Wigner’s theorem, we prove that the energy levels internal to the spaceship must undergo spontaneous discretization. The level separation turns out to be finely tuned so that, after completing a roundtrip of the curve, all systems are back to their initial state. This implies, for example, that the memories of an observer inside the spaceship are necessarily erased by the end of the journey. More in general, if there is an increase in entropy, a Poincaré cycle will eventually reverse it by the end of the loop, forcing entropy to decrease back to its initial value. We show that such decrease in entropy is in agreement with the eigenstate thermalization hypothesis. The non-existence of time-travel paradoxes follows as a rigorous corollary of our analysis.
@@ I Introduction
An insightful thermodynamic picture was proposed in [2]. Let us briefly review the main idea. We start by noticing that, according to our current understanding of physics, entropy increase is the only physical law that enables us to draw a fundamental distinction between past and future [^1] Ignoring phenomena that may depend on the chosen interpretation of quantum mechanics [3, 4]. This gives us a local criterion to determine the direction of time. In particular, if a thermally isolated system (see [5, §11] for an operational definition) occupies an event $x$ in spacetime, the “future” lightcone at $x$ is the one that points in the direction of increasing entropy of such thermodynamic system. Let us apply this principle to a CTC, assuming that the spaceship can indeed be approximated as a thermally isolated system.
Since a CTC is a compact set, and the entropy $S$ is a continuous function, there exists an event $x_{0}$ in the curve where the entropy of the spaceship is minimal, and another event $x_{f}$ where the entropy is maximal. Let us then break the closed curve into two open segments, which begin in $x_{0}$ and end in $x_{f}$. Since $S(x_{0})\leq S(x_{f})$, the entropy grows (on average) from $x_{0}$ to $x_{f}$ on both segments. Hence, the thermodynamic time begins in $x_{0}$ and ends in $x_{f}$ on both branches of the CTC, see Figure 1. Based on this fact, it is argued in [2] that, if an observer begins their journey in $x_{0}$, and travels counterclockwise on the green line in Figure 1, then, after they cross the event $x_{f}$, we are no longer allowed to say that their subjective time “keeps growing”. Instead, all thermodynamic processes (including biological processes such as memory formation and aging) are reversed. For this reason, strictly speaking, the story of this observer ends in $x_{f}$. The blue line contains an alternative story, where another version of the observer travels from $x_{0}$ to $x_{f}$ clockwise. In conclusion: A closed timelike curve is not a loop in time, but two parallel timelines, both of which begin and end in the same events (namely $x_{0}$ and $x_{f}$).
@@
The purpose of the present article is to show, using quantum statistical mechanics, that the qualitative picture of [2] outlined above is “correct”, in the sense that it agrees with our current understanding of the laws of physics. To achieve this goal, we study the evolution of a “spaceship” (i.e., of a closed quantum system) traveling along a CTC in a Gödel-type universe [6, 7]. To be as rigorous as possible, we will ground all our claims on commonly accepted quantum mechanical principles and theorems, without invoking any assumption that may be regarded as controversial. We will verify that the thermodynamic picture of Figure 1 emerges naturally.
@@
The only way out of these paradoxes is to appeal to a self-consistency principle, according to which the only histories of the Universe that actually happen are those where everything works out consistently [8, 9]. For example, in Alice’s experiment above, it may be assumed that the coin must always give consistent outcomes on both parallel histories. In a classical world, this must be enforced as an additional constraint on the Cauchy problem [10, 11], while in a quantum world self-consistency automatically holds as an intrinsic feature of the evolution operator [11, 12, 13, 14]. Indeed, we will show in section II.3 that, in our spacetimes of interest, self-consistency of histories follows as a corollary of Wigner’s theorem [15], and it holds for any initial quantum state |$\ket{\Psi}$. This is crucial for us, since it would be hard to make statistical mechanical claims for a system with constraints that are non-local in time.
@@ II Quantum Mechanics in a Gödel-type universe
In this section, we show that the energy levels of a closed system traveling along a CTC are quantized, and forced to be multiples of a fundamental discrete unit, whose value is inversely proportional to the length of the curve. Since the present article is not concerned with the backreaction of the system on the spacetime geometry, we will treat the spacetime metric as a fixed classical background, and assume that the quantum system has a very small mass.
@@ II.1 Brief description of the metric
A Gödel-type universe [6, 7] is a rotating universe that admits CTCs. We work in a cylindrical coordinate system $\{t,\phi,r,z\}$, where $t\in\mathbb{R}$ is a time coordinate, $\phi=[0,2\pi]$ is the cylindrical angle (with $0\equiv 2\pi$), $r\geq 0$ is the distance from the cylindrical axis, and $z\in\mathbb{R}$ parametrizes the axis. For the physics we are interested in, it will suffice to postulate the following general metric:
@@
where $A$, $B$, $C$ are generic functions of $r$. Gödel’s metric can be expressed in the above form [16, §5.7], with
$\begin{split}A={}&-1\,,\\ B={}&\sqrt{2}\sinh^{2}r\,,\\ C={}&\sinh^{2}r-\sinh^{4}r\,.\\ \end{split}$ (2)
Closed timelike curves are guaranteed to exist whenever $C(r_{0})<0$ for some $r_{0}$. In fact, when this happens, $g(\partial_{\phi},\partial_{\phi})<0$, meaning that $\partial_{\phi}$ is timelike at $r_{0}$. Its integral curves are therefore timelike, and have the form
@@ II.2 Symmetries of the metric and their quantum realizations
The Gödel metric is known to possess 5 linearly independent Killing vectors [17]. For the purposes of our work, only one is relevant, namely $\partial_{\phi}$, which is the generator of rotations around the $z$ axis in all universes of the form (1). This family of transformations is an abelian one-parameter group of symmetries of the spacetime, which act as follows:
@@
If we treat the spacetime as a fixed background, the laws of physics must be invariant under all transformations $\mathcal{G}_{b}$. According to Wigner’s theorem, if a quantum theory is invariant under a group of symmetries, its Hilbert space is the carrier space of a unitary representation of such group [15, §2.2]. For the group (4), the corresponding unitary representation takes the form (by Stone’s theorem [18])
@@
A key feature of the group $\mathcal{G}_{b}$ is that, due to the identification $0\equiv 2\pi$ for the variable $\phi$, we have the relation $\mathcal{G}_{2\pi}=1$. The group (5), being a representation of (4), must reflect this constraint. In particular, the action of the operator $U(2\pi)$ on a quantum state |$\ket{\Psi}$ must return the same state |$\ket{\Psi}$, up to a constant phase $e^{i\alpha}$ independent of |$\ket{\Psi}$ [15] (if there are superselection rules, we must work within a coherent space with constant $\alpha$ [19, §1-2]). Therefore, we have the following operatorial identity:
@@
[…]
@@ II.3 Quantum mechanics on a CTC
This implies that, if the spaceship is thermally isolated, it evolves in proper time $\tau$ according to the unitary operators $e^{iJ\tau/\sqrt{-C(r_{0})}}$, up to an arbitrary phase. The formal proof is provided in Appendix A. Thus, the Hamiltonian operator generating the spaceship’s evolution is
@@ II.4 Spontaneous discretization of the energy levels
II.5 Comment of finite-size effects
In the above derivation, we treated the spaceship as a point-like system. However, in reality, the spaceship has a finite size, and the revolution period may be different on opposite edges of the spaceship. When this effect is relevant, all the above equations still hold. We only need to interpret $r_{0}$ as the location of an idealized clock inside the spaceship, and define $\mathcal{T}$ there. Then, our analysis above still applies. The only subtlety is that we need to revise our interpretation of $\tau$. In particular, if a physical process in the spaceship lasts a time $\Delta\tau$, this does not coincide with the proper time experienced by the particles that participate in the process. Instead, it is the duration of the process as perceived by an observer located at $r_{0}$. The proof is provided in Appendix B. For our purposes, the present distinction is irrelevant.
@@ III Spontaneous inversion of the entropic arrow of time
In this section, we show that the thermodynamic picture illustrated in figure 1 (and introduced in [2]) is valid (at least in Gödel-type universes), and follows directly from equation (10). In the following, we will model the spaceship as an idealized box with perfectly reflecting walls. This is necessary, because the second law of thermodynamics applies only to thermally isolated systems, to which we can assign a Hamiltonian [5, §11].
@@ III.1 A simple example: Spontaneous recombination of an unstable particle
According to equation (10), the Hamiltonian of the system takes the general form
@@
where $P_{n}=P_{n}^{\dagger}$ are orthogonal projectors, such that $P_{n}P_{m}=\delta_{nm}P_{n}$ and $\sum_{n\in\mathbb{Z}}P_{n}=1$ [22, Ch.1, §6.9]. The probability amplitude of observing again the particle $\Psi$ after a proper time $\tau$ has passed [23] is given by
@@ III.3 Discussion
Both plots 2 and 3 confirm the qualitative picture outlined in [2], and illustrated in Figure 1. In fact, if we identify the instant $\tau=0$ with the minimum-entropy event $x_{0}$, the entropy grows in both time directions within a finite neighborhood of $x_{0}$ (see [26] for quantitative claims). However, as we move far from $x_{0}$, at some point we meet an event $x_{f}$ where the entropy is maximal, and the second law of thermodynamics begins to turn back. In Figure 2, the event $x_{f}$ occurs at $\tau=\mathcal{T}/2$. In fact, around $\mathcal{T}/2$, the probability $\mathcal{P}_{\Psi}$ is closest to 00, which is the thermodynamic equilibrium value, see equation (11). In Figure 3, the interior of the spaceship is in thermodynamic equilibrium across the interval $\tau/\mathcal{T}\in[0.05,0.95]$. Thus, there is no preferred direction of time in this historical age, and we can pick $x_{f}$ at will.
@@ IV.1 Clocks
It was noted in [2] that, on a CTC, the ticking frequency of all “good” clocks must be an integer multiple of $2\pi/\mathcal{T}$. In fact, self-consistency demands that, at the end of the trip, the hand of the clock must be back where it started. This leads one to the following logical implication: If the length $\mathcal{T}$ of the CTC is a non-integer number of seconds, a second-based clock traveling on the spaceship must suffer from inherent imperfections. These imperfections have a microscopic origin and arise due to quantum effects. Let us show this with a quick example.
@@ IV.2 Spontaneous memory erasure
Memory is the result of an interaction, where an object leaves traces of its initial state in the later state of a “memory-keeper” [3], which may be a measurement device or a living being. Let us assume, for clarity, that the state space of the spaceship can be (approximately) decomposed as the tensor product between the object’s state space and the memory-keeper’s state space. Then, given an orthonormal family of object’s states |$\ket{m}$, which the memory keeper can distinguish through senses, memory collection can be modeled as the following process:
@@ IV.3 The minimum-entropy event
Note that none of the above constitutes a real paradox. Complex structures (like brains) fluctuate into existence in all ergodic systems. The difference is in the waiting time $\tau_{W}$ before these structure forms. On CTCs, $\tau_{W}{\equiv}\mathcal{T}$. In an ergodic system with $N$ particles, $\tau_{W}\,{\propto}\,e^{N}$ [20, §4.5].
IV.4 Can I meet myself from the future?
One of the most confusing issues surrounding time travel is the possibility that an observer might interact with a version of themselves coming from the future. It is often suggested that this eventuality may give rise to retro-causality paradoxes. Let us see what quantum statistical mechanics has to say about this eventuality.

### v2

@@ Abstract
We study the internal dynamics of a hypothetical spaceship traveling on a close timelike curve in an axially symmetric universe. We choose the curve so that the generator of evolution in proper time is the angular momentum. Using Wigner’s theorem, we prove that the energy levels internal to the spaceship must undergo spontaneous discretization. The level separation turns out to be finely tuned so that, after completing a roundtrip of the curve, all systems are back to their initial state. This implies, for example, that the memories of an observer inside the spaceship are necessarily erased by the end of the journey. More in general, if there is an increase in entropy, a Poincaré cycle will eventually reverse it by the end of the loop, forcing entropy to decrease back to its initial value. We show that such decrease in entropy is in agreement with the eigenstate thermalization hypothesis. The non-existence of time-travel paradoxes follows as a rigorous corollary of our analysis.
@@ I Introduction
An insightful thermodynamic picture was proposed in [2]. Let us briefly review the main idea. We start by noticing that, according to our current understanding of physics, entropy increase is the only physical law [^1] Ignoring phenomena that may depend on the chosen interpretation of quantum mechanics. that enables us to draw a fundamental distinction between past and future [3, 4]. This gives us a local criterion [^2] Here, “local” means that the criterion applies to regions of spacetime that are small compared to the spacetime geometry, but still large enough to host a closed thermodynamic system. to determine the direction of time. Namely, if a thermally isolated system (see [5, §11] for an operational definition) occupies an event $x$ in spacetime, the “future” lightcone at $x$ is the one that points in the direction of increasing entropy of such thermodynamic system. Let us apply this principle to a CTC, assuming that the spaceship can indeed be approximated as a thermally isolated system.
@@
Since a CTC is a compact set, and the entropy $S$ is a continuous function, there exists an event $x_{0}$ in the curve where the entropy of the spaceship is minimal, and another event $x_{f}$ where it is maximal [^3] To keep the discussion simple, we assume that $x_{0}$ and $x_{f}$ are unique. The generalization to more complicated scenarios with multiple extrema is straightforward. The case where the entropy is constant along the whole CTC is not interesting for us, since this article focuses on non-equilibrium processes on CTCs.. Let us then break the closed curve into two open segments, which begin in $x_{0}$ and end in $x_{f}$. Since $S(x_{0})\leq S(x_{f})$, the entropy grows (on average) from $x_{0}$ to $x_{f}$ on both segments. Hence, the thermodynamic time begins in $x_{0}$ and ends in $x_{f}$ on both branches of the CTC, see Figure 1. Based on this fact, it is argued in [2] that, if an observer begins their journey in $x_{0}$, and travels counterclockwise on the green line in Figure 1, then, after they cross the event $x_{f}$, we are no longer allowed to say that their subjective time “keeps growing”. Instead, all thermodynamic processes (including biological processes such as memory formation and aging) are reversed. For this reason, strictly speaking, the story of this observer ends in $x_{f}$. The blue line contains an alternative story, where another (time-reversed) version of the observer travels from $x_{0}$ to $x_{f}$ clockwise. In conclusion: From the perspective of an observer traveling on the CTC, a closed timelike curve is not a loop in (entropic) time. Rather, it is two parallel entropic timelines, both of which begin and end in the same events (namely $x_{0}$ and $x_{f}$).
The purpose of the present article is to show, using quantum statistical mechanics, that the qualitative picture of [2] outlined above is “correct”, in the sense that it agrees with our current understanding of the laws of physics. To achieve this goal, we study the evolution of a “spaceship” (i.e., of a closed quantum system) traveling along a CTC that is the orbit of a local Killing vector. We will use, as our reference spacetimes, the Gödel-type universes [6, 7], since they have a very simple geometry. However, our main conclusions apply to a wide variety of spacetimes, notably Kerr and [8, 9, 10, 11, 12]. To be as rigorous as possible, we will ground all our claims on commonly accepted quantum mechanical principles and theorems, without invoking any assumption that may be regarded as controversial. We will verify that the thermodynamic picture of Figure 1 emerges naturally.
@@
The only way out of these paradoxes is to appeal to a self-consistency principle, according to which the only histories of the Universe that actually happen are those where everything works out consistently [13, 14]. For example, in Alice’s experiment above, it may be assumed that the coin must always give consistent outcomes on both parallel histories. In a classical world, this must be enforced as an additional constraint on the Cauchy problem [15, 16, 17, 18], while in a quantum world self-consistency automatically holds as an intrinsic feature of the evolution operator [18, 19, 20, 21]. Indeed, we will show in section II.3 that, in our spacetimes of interest, self-consistency of histories follows as a corollary of Wigner’s theorem [22], and it holds for any initial quantum state |$\ket{\Psi}$. This is crucial for us, since it would be hard to make statistical mechanical claims for a system with constraints that are non-local in time.
@@ II Quantum Mechanics in a Gödel-type universe
In this section, we show that the energy levels of a closed system traveling along a CTC are quantized, being multiples of a fundamental discrete unit, whose value is inversely proportional to the length of the curve. Since the present article is not concerned with the backreaction of the system on the spacetime geometry, we treat the spacetime metric as a fixed classical background, and assume that the quantum system has very small mass.
@@ II.1 Brief description of the metric
A Gödel-type universe [6, 7, 23] is a rotating universe that admits CTCs. We work in a cylindrical coordinate system $\{t,\phi,r,z\}$, where $t\in\mathbb{R}$ is a time coordinate, $\phi=[0,2\pi]$ is the cylindrical angle (with $0\equiv 2\pi$), $r\geq 0$ is the distance from the cylindrical axis, and $z\in\mathbb{R}$ parametrizes the axis. For the physics we are interested in, it will suffice to postulate the following general metric:
@@
where $A$, $B$, $C$ are generic functions of $r$. Gödel’s metric can be expressed in the above form [24, §5.7], with
$A{=}{-}1\,,\quad B{=}\sqrt{2}\sinh^{2}r\,,\quad C{=}\sinh^{2}r{-}\sinh^{4}r\,.$ (2)
Closed timelike curves are guaranteed to exist whenever $C(r_{0})<0$ for some $r_{0}$ [23]. In fact, when this happens, $g(\partial_{\phi},\partial_{\phi})<0$, meaning that $\partial_{\phi}$ is timelike at $r_{0}$. Its integral curves are therefore timelike, and have the form
@@ II.2 Symmetries of the metric and their quantum realizations
The Gödel metric is known to possess 5 linearly independent Killing vectors [25]. For the purposes of our work, only one is relevant, namely $\partial_{\phi}$, which is the generator of rotations around the $z$ axis in all Gödel-type universes of the form (1). This family of transformations is an abelian one-parameter group of symmetries of the spacetime, which act as follows:
@@
If we treat the spacetime as a fixed background, the laws of physics must be invariant under all transformations $\mathcal{G}_{b}$. According to Wigner’s theorem, if a quantum theory is invariant under a group of symmetries, its Hilbert space is the carrier space of a unitary representation of such group [22, §2.2]. For the group (4), the corresponding unitary representation takes the form (by Stone’s theorem [26])
@@
A key feature of the group $\mathcal{G}_{b}$ is that, due to the identification $0\equiv 2\pi$ for the variable $\phi$, we have the relation $\mathcal{G}_{2\pi}=1$. The group (5), being a representation of (4), must reflect this constraint. In particular, the action of the operator $U(2\pi)$ on a quantum state |$\ket{\Psi}$ must return the same state |$\ket{\Psi}$, up to a constant phase $e^{i\alpha}$ independent of |$\ket{\Psi}$ [22] (if there are superselection rules, we must work within a coherent space with constant $\alpha$ [27, §1-2]). Therefore, we have the following operatorial identity:
@@
A more detailed proof of (6), which makes explicit use of the definition of symmetry in quantum mechanics, is provided in Appendix A.
@@ II.3 Quantum mechanics on a CTC
This implies that, if the spaceship is thermally isolated, it evolves in proper time $\tau$ according to the unitary operators $e^{iJ\tau/\sqrt{-C(r_{0})}}$, up to an arbitrary phase. The formal proof is provided in Appendix B. Thus, the Hamiltonian operator generating the spaceship’s evolution is
@@ II.4 Spontaneous discretization of the energy levels
II.5 Comment on finite-size effects
In the above derivation, we treated the spaceship as a point-like system. However, in reality, the spaceship has a finite size, and the revolution period may be different on opposite edges of the spaceship. When this effect is relevant, all the above equations still hold, but we need to reinterpret $r_{0}$ as the exact location of a clock somewhere in the spaceship, and define $\mathcal{T}$ there. Then, our analysis above still applies (see Appendix C.1 for the proof). The only subtlety now is that we need to revise our interpretation of $\tau$. In particular, if a physical process in the spaceship lasts a time $\Delta\tau$, this does not coincide with the proper time experienced by the particles that participate in the process. Instead, it is the duration of the process as perceived by an observer located in the proximity of the clock (at $r_{0}$). The proof is provided in Appendix C.2. For our purposes, the present distinction is irrelevant.
@@ III Spontaneous inversion of the entropic arrow of time
In this section, we show that the thermodynamic picture illustrated in figure 1 (and introduced in [2]) is valid (at least in axisymmetric universes), and follows directly from equation (10). In the following, we will model the spaceship as an idealized box with perfectly reflecting walls. This is necessary, because the second law of thermodynamics applies only to thermally isolated systems, to which we can assign a Hamiltonian [5, §11].
@@ III.1 A simple example: Spontaneous recombination of an unstable particle
According to equation (10), the Hamiltonian of the system takes the general form [26, §3.1, Theorem 3.7]
@@
where $P_{n}=P_{n}^{\dagger}$ are orthogonal projectors, such that $P_{n}P_{m}=\delta_{nm}P_{n}$ and $\sum_{n\in\mathbb{Z}}P_{n}=1$ [30, Ch.1, §6.9]. Note that the spectrum of $H$ does not need to be the whole set $2\pi\mathbb{Z}/\mathcal{T}$, because some projectors $P_{n}$ may vanish. The probability amplitude of observing again the particle $\Psi$ after a proper time $\tau$ has passed [31] is given by
@@ III.3 Discussion
Both plots 2 and 3 confirm the qualitative picture outlined in [2], and illustrated in Figure 1. In fact, we recall that, if the non-equilibrium deviation is not too large, the entropy of a macrostate with a given value of $A$ can be expanded as follows [5, §110]:
$S(A)=S(\bar{A})-\dfrac{1}{2}|S^{\prime\prime}(\bar{A})|\,(A{-}\bar{A})^{2}.$ (19)
Hence, we can identify the instant $\tau=0$ with the minimum-entropy event $x_{0}$, and the entropy grows in both time directions within a finite neighborhood of $x_{0}$ (see [34] for quantitative claims). However, as we move far from $x_{0}$, at some point we meet an event $x_{f}$ where the entropy is maximal, and the second law of thermodynamics begins to turn back. In Figure 2, the event $x_{f}$ occurs at $\tau=\mathcal{T}/2$. In fact, around $\mathcal{T}/2$, the probability $\mathcal{P}_{\Psi}$ is closest to 00, which is the thermodynamic equilibrium value, see equation (11). In Figure 3, the interior of the spaceship is in thermodynamic equilibrium across the interval $\tau/\mathcal{T}\in[0.05,0.95]$. Thus, there is no preferred direction of time in this historical age, and we can pick $x_{f}$ at will.
Note that the qualitative behavior in Figure 3 is essentially universal to all statistical systems fulfilling the eigenstate thermalization hypothesis and such that (9) holds. In fact, all such systems must (a) evolve back to their initial state at time $\tau=\mathcal{T}$, and (b) spend most of their time in the most likely state, which is $\langle A\rangle=\bar{A}$.
@@ IV.1 Clocks
It was noted in [2] that, on a CTC, the ticking frequency of all “good” clocks must be an integer multiple of $2\pi/\mathcal{T}$. In fact, self-consistency demands that, at the end of the trip, the hand of the clock must be back where it started. This leads one to the following logical implication: If the CTC’s period $\mathcal{T}$ lasts a non-integer number of seconds, all second-based clocks traveling on the CTC must suffer from “production defects”, since they will tick an integer number of times over a non-integer number of seconds. These defects have a microscopic origin and arise due to quantum effects. Let us show this with a quick example.
@@ IV.2 Spontaneous memory erasure
Memory may be schematically modeled as the result of an interaction, where an object leaves traces of its initial state in the later state of a “memory-keeper” [3], which may be a measurement device or a living being. Let us assume, for clarity, that the state space of the spaceship can be (approximately) decomposed as the tensor product between the object’s state space and the memory-keeper’s state space. Then, given an orthonormal family of object’s states |$\ket{m}$, which the memory-keeper can distinguish through senses, memory collection can be schematically modeled as the following process [^4] Of course, this is only a simplified picture, since the problem of how observables are measured and memory is kept is more delicate than what equations (21) and (22) may suggest. However, the main result should hold in general.:
@@ IV.3 The minimum-entropy event
Note that none of the above constitutes a real paradox. Complex structures (like brains) fluctuate into existence in all ergodic systems. The difference is in the waiting time $\tau_{W}$ before these structure forms. On CTCs, $\tau_{W}{\equiv}\mathcal{T}$. In an ergodic system with $N$ particles, $\tau_{W}\,{\sim}\,e^{N}$ [28, §4.5].
IV.4 Speculation: Can I meet myself from the future?
One of the most confusing issues surrounding time travel is the possibility that an observer might interact with a version of themselves coming from the future. It is often suggested that this eventuality may give rise to retro-causality paradoxes. Let us see what quantum statistical mechanics has to say about this eventuality. We must warn the reader that the following discussion is rather speculative. Yet, we believe that it is worth pondering.